# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v4

@@ Abstract
We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear model for the given problem and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments and covers all the requirements posed by the operator of the network. The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. At the same time, the solutions that were obtained were feasible. Moreover, though they are probabilistic (heuristics) they offer a valid alternative by returning a range of possible solutions the dispatcher can choose from. And, most importantly, they outperform classical solvers in some cases.
@@
railway rescheduling, conflict management, heterogeneous urban railway network, quantum annealing, hybrid quantum-classical heuristics
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems. Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays (Ge et al., 2022). Examples of such disturbances include: late train departures/arrivals, extended dwell times. These can last from several minutes up to hours. A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too (Cacchiani et al., 2014). In the present work we focus on rescheduling of trains, which is the network operator’s perspective, even though we also consider (partial) track closures, which are are typically considered as disruptions. In our examples, however they will be managed by modifying existing train paths. The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f. Törnquist (2007) and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
@@
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays (Lange and Werner, 2018), the maximal delay that cannot be avoided (D’Ariano et al., 2007), or fuel consumption measures  (Harrod, 2011). Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time. The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints (Szpigel, 1973; Mascis and Pacciarelli, 2002). It has an extensive literature, already summarized in a number of reviews, including those by Cacchiani et al. (2014), by Lamorgese et al. (2018), or by Philip and Swapnesh (2022).
@@
The maybe most frequently applied modeling strategy is based on alternative graphs (D’Ariano et al., 2007; Mascis and Pacciarelli, 2002) which can be formulated as as a mixed integer linear program, often referred to as the big-M model. As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and the combinations of these; see the work of Lamorgese et al. (2018) for a review of the model and the methods.
@@
An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling (Caimi et al., 2012; Lusby et al., 2013; Meng and Zhou, 2014). As this approach results in a $0-1$ program, which is promising when one considers the use of $0-1$ solvers such as quantum hardware. Indeed, job-shop problems addressed on pure quantum hardware (Venturelli et al., 2015), and even in certain hybrid approaches (Kurowski et al., 2020) use similar models, and our previous work (Domino et al., 2023) which was the first to apply a quantum solver to railway dispatching problems also followed this route. In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP). We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances. In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
@@
In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model by Domino et al. (2022b). Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
@@
We recognize several important scientific gaps. First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances. Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling. In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks.
@@
The paper is organized as follows. In Section 2 we briefly review state-of-the-art of quantum annealing and its applications. In Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
@@ 2 Quantum annealing
Quantum computing devices became available for practical computational purposes in the last few years. Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics (Ising, 1925; Bian et al., 2010). The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
Hence, a QUBO objective can be expressed as the energy of the Ising system: the configuration of the pairwise couplings of the spins and the applied local fields, which are tunable in quantum hardware. According to the adiabatic theorem of quantum mechanics (Avron and Elgart, 1999), tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out. Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems. Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  (Das and Chakrabarti, 2008).
In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs. They output a number of configurations typically close to, and optionally including the optimal one.
@@
The class of QUBO problems is well studied and have a number of applications; we refer the reader to a recent monograph edited by Punnen (2022). Linear unconstrained binary problems can also be considered as a special case of QUBOs, while linear and quadratic constraints can be taken into account with penalties (Gusmeroli et al., 2022). Using appropriate binary encoding of the variables, integer programming problems can also be treated. Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
@@
The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled. The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system. This embedding is a hard problem itself (Zbinden et al., 2020). It often requires to couple multiple physical qubits to represent a logical bit of the problem. As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem are encoded with a limited accuracy.
@@
To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently. Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a), including the one used in the present work, belong to this family. Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
@@
Currently, the applicability of QA technologies is being explored. In the railway context, the first applications of QA can be found in train dispatching/rescheduling (Domino et al., 2022b) and rolling stock planning (Bickert et al., 2021). In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us (Domino et al., 2023). A follow-up paper (Domino et al., 2022b) laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research, which is gaining increasing attention, e.g. the recent work of Xu et al. (2023).
@@ 3 Problem description
We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations. Also, we consider an urban railway network with mixed tracks from single, double, to quadruple tracks.
@@
We model a railway network with edges and nodes of a graph. Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines). Each edge is composed of one or more tracks. Each track consists of block sections defined by pairs of signals (Lusby et al., 2013). Each block section can be occupied by at most one train at a time. A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains. We assume a 2-block signaling system, meaning that two free blocks are required between the consecutive trains. We also assume a green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
@@
Each node is composed of station tracks (blocks) and interlocking areas. One station track can be occupied by one train at a time. Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at the time.
@@
A train’s route is the sequence of blocks the train passes during its journey. A train path is a sequence of arrival and departure times of a particular train assigned to a train route. We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time, or passing time, resulting in integer variables. This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables. The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains, encoded later on precedence binary variables.
@@
In the network, two trains can follow each other, i.e. keep the given order, meet and pass (M-P) when going the opposite directions, and meet and overtake (M-O), i.e. change the order, when going the same direction. Single-track lines are designed for bidirectional traffic, double track lines for unidirectional, one for each direction, and multi-track lines can be a combination of unidirectional and bidirectional. On single-track lines, M-Ps and M-Os are only possible at stations. To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single-track line heading in opposite directions.
@@
Trains in the same direction preserve their order between stations, and keep the minimal headway time between each other. To prescribe these we determine set $\mathcal{J}^{2}_{\text{headway}}$: the set of all pairs of trains that can potentially violate the minimal headway condition.
@@
The train traffic is scheduled based on the given timetable. The timetable contains all the train paths. In the train rescheduling problem, we are given an initial timetable that is conflict-free. However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc. Following Corman et al. (2012), a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously. We assume that the conflict may occur either on the railway line or at the station. Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time. Conflicts at stations include claiming a station track by two trains at the same time, or claiming a station switch (in the interlocking area) by two trains at the same time.
@@
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. We will minimize a function of these secondary delays.
An important element of our approach is that from among all stations, we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only. Hence, as decision stations, we select those stations, where routes of trains intersect, where trains start or terminate, or where the selected part of the network is bounded. The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths have to be made with respect to these stations. (If we change some of the trains’ routes in a re-routing process, the new model with new parameters is developed, e.g. additional decision stations may appear.) In what follows, by a station we always mean a decision station. This also means that non-decision stations appear only through the parameter values in the model, they do not appear as indices of decision variables or parameters. Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without a detailed layout. We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot. The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.
@@
In this way our model shares features of both microscopic and macroscopic models. At the decision stations it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting. Meanwhile from the point of view of the rest of the network it is a macroscopic model.

### v5

@@ Abstract
We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments. All the requirements posed by the operator of the network are modeled. The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@
[…]
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policies, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems. Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [18]. Examples of such disturbances include late train departures/arrivals, extended dwell times, etc. Their duration can be from several minutes up to hours. A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too [5]. In the present work, we focus on train rescheduling, which is the network operator’s perspective. Even though we also consider (partial) track closures, which are typically considered as disruptions, in our examples they will be managed by modifying existing train paths. The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f. [37] and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
@@
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for the efficient suppression of the consequences of disturbances. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22]. Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost in real time. The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints [36, 29]. It has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
@@
The maybe most frequently applied modeling strategy is based on alternative graphs [10, 29], resulting in a disjunctive program. Such a model can also be converted into a mixed integer linear program referred to as the big-M model. As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
@@
An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though the so-arising $0-1$ programs can be large, this method has been applied for both timetabling and dispatching/rescheduling [6, 28, 30]. Meanwhile, $0-1$ programs are suitable as input for certain quantum hardware types. Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models, and our previous work [15] which was the first to apply a quantum solver to railway dispatching problems also followed this route. In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP). We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances. In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
@@
In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We do not include train cancellations in the model. We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model by [17]. Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
@@
We recognize several important scientific gaps. First, existing optimization models do not scale well: they can become huge and inefficient for larger real-life instances. Exploiting the characteristics of a given railway scenario, one can possibly construct simplified models whose size remains small enough to be addressed with a suitable exact or heuristic solver, yet they capture the relevant optimization objectives and constraints. For the particular railway scenario and solver studied in the present paper, no such model is known. Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling. In this paper, we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of an urban railway network.
@@
The paper is organized as follows. In Section 2 we briefly review the state of the art of quantum annealing and its applications. In Section 3 we describe the problem under investigation, in Section 4 we present our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results. In Section 7 conclusions are drawn.
@@ 2 Quantum annealing
Quantum computing devices became available for practical computational purposes in the last few years. Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics [23, 3]. The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
@@
The class of QUBO problems is well studied and has a number of applications; we refer the reader to a recent monograph edited by [32]. Linear binary problems can also transformed to QUBOs; linear and quadratic constraints can be taken into account with penalties [21]. Using appropriate binary encoding of the variables, integer programming problems can also be treated. Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
Expressing a QUBO as the energy of an Ising system one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which can be realized in tunable quantum hardware. According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out. Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems. Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  [11].
In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs. They output a number of configurations typically close to, and possibly including the optimal one.
@@
The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled. The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system. This embedding is a hard problem itself [41]. It often requires to couple multiple physical qubits to represent a logical bit of the problem. As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem can only be set with limited accuracy.
@@
To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine to address hard problem instances more efficiently. Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ [8], including the one used in the present work, belong to this family. Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
@@
Currently, the applicability of QA technologies is being explored. In the railway context, the first applications of QA can be found in train dispatching/rescheduling [17] and rolling stock planning [4]. In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [15]. A follow-up paper by [17] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research, which is gaining increasing attention, e.g. the recent work of [40].
@@ 3 Problem description
We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations. The considered traffic takes place in an urban railway network with mixed track systems: segments of single, double, to quadruple tracks.
@@
We model a railway network with edges and nodes of a graph. Figure 1 depicts an example of a network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines). Each edge is composed of one or more tracks. Each track consists of block sections defined by pairs of signals [28]. Each block section can be occupied by at most one train at a time. A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains. To place the model into a real-life railway context, we adopt a 2-block signaling system, meaning that two free blocks are required between the consecutive trains. The 2-block signaling system is common across the Polish railway network, and typical for conventional primary railway lines. The signaling system affects the model through the way of calculating parameter values (headway times), hence, the model can be applied to various safety systems. We assume green way policy, that is, trains have the free way to move at the maximal allowed speed between stations. This results in constant running times for each block.
@@
Each node is composed of station tracks (blocks) and interlocking areas. The conflicts described by our model are the following. Each station track can be occupied by at most one train at a time. Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at a time.
@@
A train’s route is the sequence of blocks the train passes during its journey. A train path is a sequence of arrival and departure times of a particular train assigned to a train route. We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time, or passing time, resulting in integer variables. This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning binary variables. The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains which is modeled by the binary precedence variables.
@@
In the network, two trains can follow each other keeping their given order. They can meet and overtake (M-O) when they go in the same direction, thereby changing their order. A pair of trains heading in the opposite direction can meet and pass (M-P). On single-track lines, M-Ps and M-Os are only possible at stations. This has to be enforced by constraints in our model. To maintain a reasonable number of constraints, the set of all pairs of trains that are involved in such constraints, that is, which can potentially meet on any single-track line segment in opposite directions, can be determined in advance from the model instance data.
@@
Similarly, trains in the same direction preserve their order between stations and keep the minimal headway time between each other. This does not hold for trains running in the same direction on multiple-track segments on different tracks. The set of all pairs of trains that can potentially violate the minimal headway condition can also be predetermined, thereby minimizing the number of the respective constraints.
@@
The train traffic is scheduled based on the given timetable. The timetable contains all the train paths. In the train rescheduling problem, we are given an initial timetable that is conflict-free. However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc. The disturbances appear in our model via initial delays: the delays of trains at the beginning of their journey through the network. The initial delays have fixed values in each of our problem instances; the study of the statistics of delays because of random initial delays [35] is beyond the scope of the present study. Following [7], a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously. Conflicts can occur either on railway lines or at the stations. Possible conflicts that can occur on the lines include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time. Conflicts at stations include claiming a station track by two trains at the same time or claiming a station switch (in the interlocking area) by two trains at the same time.
@@
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in additional delays of trains, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. We will minimize the function of these secondary delays.
First we determine the subset of decision stations: these are the nodes of the network (typically: stations) where the events that are directly affected by the decisions can take place. These events are: starting, termination, entering or leaving the modeled part of the network, meet and pass, meet and overtake, and the joining and splitting of train routes.
The choice of decision stations depends on railway operation practices and actual traffic circumstances: stations where trains are allowed to meet and pass or meet and overtake, network junctions, and the boundaries of the selected part of the network have to be considered as decision stations. Meanwhile, there can be stations with an infrastructure supporting meet and pass or meet and overtake, which are not commonly used for this purpose under normal circumstances. These are normally not considered as decision stations. In the case of disruptions, however, some of them can be introduced as additional decision stations upon the network operator’s decision. From the model’s point of view, the decision stations are those to which decision variables are tied to. In what follows, we use the term "station" for decision stations only, unless otherwise stated.
The model does not include decision variables tied to non-decision stations. For instance, the precedence of trains cannot be changed on non-decision stations. They do have their impact on parameter values in the model though. Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without a detailed layout. We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot. Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
@@
Our model shares features of both microscopic and macroscopic models. At the decision stations, it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting. Meanwhile, from the point of view of the rest of the network, it is a macroscopic model.